OpenStudy (anonymous):

simplify by removing factors of 1. 14x^2-56/49x^2-196

OpenStudy (saifoo.khan):

\[\frac{14(x-2)(x+2)}{49(x-2)(x+2}\]

OpenStudy (anonymous):

14(x^2-4)/49(x^2-4) 14/49 2/7

OpenStudy (saifoo.khan):

\[\frac{14}{49}\]\[\frac{2}{9}\]